gibster44:

Jeffrey is saving up for a down payment on a car. He plans to invest $2,000 at the end of every year for 4 years. If the interest rate on the account is 2.15% compounding annually, what is the present value of the investment? (2 points) a $7,587.82 b $5,033.72 c $8,261.72 d $15,252.94

supie:

\(\text{Ok so we have to use the Compound Interest Formula:}\) \(\LARGE A = P * (1 + (\frac rn))^{(n*t)}\) Where ``` >A~total amount >P~Principal >r~annual interest rate >~compounded/year >t~time ``` So we have to know all of those, then plug in the numbers. Then solve

supie:

We're solving to find the `Total amount/A` So we don't know that yet. "He plans to invest $2,000" So `P=2000` "the interest rate on the account is 2.15% " `r=2.15` "every year for 4 years" `n=1 year` `t=4 years` So now you plug in the numbers to the compound interest formula, \(A = P * (1 + (r/n))^{(n*t)}\), then solve.